# A Learnable Wavelet Transformer for Long-Short Equity Trading and Risk-Adjusted Return Optimization

### v2

@@ V Architecture
V-A Neural Wavelet Filters
@@
V-A1 FIR Convolution Kernels and Classical Discrete–time Filters
In our neural wavelet front-end, each learnable filter is implemented as a one-dimensional finite impulse response (FIR) convolution kernel. Let $x[n]$ denote a discrete-time input signal and let $\boldsymbol{\theta}=(\theta_{0},\dots,\theta_{L-1})$ be the convolution kernel of length $L$. A standard conv1d layer with stride 1 and no dilation computes
@@ V-A1 FIR Convolution Kernels and Classical Discrete–time Filters
Thus $e^{\mathrm{j}\omega n}$ is an eigenfunction of the system and the corresponding eigenvalue
@@
V-A2 Differentiable Spectral Regularization via rFFT
Let $\boldsymbol{\theta}\in\mathbb{R}^{L}$ be a real-valued FIR kernel and $\hat{\boldsymbol{\theta}}=\mathcal{F}(\boldsymbol{\theta})$ its DFT. In matrix form, $\hat{\boldsymbol{\theta}}=\mathbf{F}\boldsymbol{\theta}$ with a fixed $\mathbf{F}\in\mathbb{C}^{L\times L}$. Hence $\mathcal{F}$ is linear and differentiable everywhere, with constant Jacobian $\mathbf{F}$. In practice we use rFFT, which is an optimized implementation of the same linear map specialized to real inputs (exploiting conjugate symmetry).
@@ V-A2 Differentiable Spectral Regularization via rFFT
$\mathcal{L}_{spec}(\boldsymbol{\theta})=\sum_{k}\rho\!\left(|\hat{\theta}_{k}|\right),$ (12)
where $\rho(\cdot)$ is differentiable. By the chain rule,
$\frac{\partial\mathcal{L}_{spec}}{\partial\boldsymbol{\theta}}=\mathbf{F}^{\mathsf{H}}\,\frac{\partial\mathcal{L}_{spec}}{\partial\hat{\boldsymbol{\theta}}}$,
with $\mathbf{F}^{\mathsf{H}}$ the Hermitian transpose (i.e., inverse DFT up to scaling). Equivalently, back-propagation through rFFT amounts to applying an inverse FFT to the frequency-domain gradient. Therefore, adding Eq. (12) on rFFT-transformed learnable kernels preserves differentiability and allows end-to-end training with standard optimizers.
To enforce complementary low/high-frequency coverage, we penalize high-frequency energy in the low-pass filter and low-frequency energy in the high-pass filter, add an overlap term to sharpen separation, and use Parseval and shape losses to constrain energy and frequency-response form; details are in Section VI-A7.
V-A3 Low-guided High-frequency Injection
@@
where $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{k}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ are projection matrices and $W_{O}$ is the output projection (multi-head variants follow the standard formulation). The final fused representation is given by a gated residual injection:
@@
VI Training Design
Training choices, such as loss, optimizer, and learning-rate schedule, strongly influence trading outcomes. Because convenient objectives may misalign with deployment metrics (return, Sharpe, draw-down), low training loss can still generalize poorly. We therefore present: (i) a differentiable surrogate objective (soft labels from future log returns + Sharpe regularization), (ii) a stabilized optimization setup for heavy-tailed financial noise, and (iii) validation/checkpoint selection based on trading performance.
VI-A Loss Function Design
@@
VI-A1 Drawbacks of Regression Losses
Many financial time-series models train neural networks as return forecasters by minimizing point-wise regression losses such as MSE or MAE. Given a future log return $\hat{\ell}_{t}$, they learn $\ell_{t}=f_{\theta}(x_{t})$ via
$\mathcal{L}_{reg}=\mathbb{E}\big[loss(\ell_{t},\hat{\ell}_{t})\big],\qquad loss\in\{\text{MSE},\ \text{MAE}\}.$ (14)
While statistically convenient, $\mathcal{L}_{reg}$ is misaligned with trading: performance is driven by the position $\tilde{w}_{t}$ and the resulting P&L $R_{t}=\tilde{w}_{t}\,\ell_{t}$, not by the numeric forecast error of $\ell_{t}$. Consequently, lower regression error does not necessarily translate to higher ROI/Sharpe. The mismatch is exacerbated by the asymmetric economic cost as a wrong sign under high exposure is far more damaging than a small magnitude error on a correct trade, whereas MSE/MAE penalize deviations symmetrically.
Moreover, under noisy and non-stationary returns, minimizing $\mathcal{L}_{reg}$ mainly encourages approximating $\mathbb{E}[\ell_{t}\mid x_{t}]$, which offers no guarantee of improved $R_{t}$ or risk-adjusted performance. Therefore, we do not treat the network as a pure forecaster. Instead, we interpret its output as a continuous trading position $\tilde{w}_{t}\in[-1,1]$ (flat if $\tilde{w}_{t}\in[-0.01,0.01]$) and optimize a trading-aligned objective,
$\mathcal{L}_{trade}=-\,\mathbb{E}\big[\tilde{w}_{t}y_{t}\big]+\lambda\,\Omega(\tilde{w}_{t}),$
@@ VI-A1 Drawbacks of Regression Losses
VI-A2 Stock Direction and Tanh Position Loss
@@
VI-A3 Sigmoid Position Loss
@@
To emphasize economically significant moves, we weight each sample by the absolute log return $|\ell_{t}|$ (e.g., $100\%$ and $50\%$ have equal $|\ell_{t}|$). The resulting Stock Sigmoid loss is
@@
VI-A4 Soft-Label Position Loss
@@
We therefore replace hard labels with probabilistic soft targets that reflect directional confidence:
$y_{t}=\sigma(k\,\ell_{t}),$
where $k>0$ controls how quickly $y_{t}$ departs from $0.5$ as $|\ell_{t}|$ increases. We choose $k$ by calibration (e.g., enforcing $\sigma(k\,\ell_{t}^{(+5\%)})\approx 0.9$), which yields $k\approx 45$ in our setting; we fix $k=45$ for all experiments across industries for reproducibility. Given $y_{t}$ and $P_{t}$, we use the soft-label cross-entropy
@@
At inference time, we map the model output to a continuous position in [1,1][-1,1] using Eq. (9), where the sign encodes long/short/flat and the magnitude controls position size. Overall, soft labels provide smoother gradients, reduce over-emphasis on extreme returns, and align the learned signal with probabilistic directional confidence.
@@ VI-A5 Overfitting Penalty
We cap the annualized ROI at $R_{\mathrm{ann}}^{\max}=1.0$ and treat it as a tunable risk-control parameter. With hourly data, $252$ trading days/year and $6.5$ hours/day, we have $H_{\mathrm{year}}=1638$. Assuming constant per-hour log return, the implied maximum batch log return is
$L_{\mathcal{B}}^{\max}=\log(1+R_{\mathrm{ann}}^{\max})\,\frac{H_{\mathcal{B}}}{H_{\mathrm{year}}},$
which yields the batch ROI threshold
$T_{\mathcal{B}}=\exp(L_{\mathcal{B}}^{\max})-1=(1+R_{\mathrm{ann}}^{\max})^{H_{\mathcal{B}}/H_{\mathrm{year}}}-1.$
We then penalize excess ROI using a one-sided quadratic hinge:
@@
where $\lambda_{\mathrm{roi}}$ controls the penalty strength.
VI-A6 Sharpe Regularizer
@@
where $r_{f}$ is the risk-free rate. We omit $r_{f}$ and use
$S=\frac{\mathbb{E}[R_{p}]}{\sigma(R_{p})},$
with $\mathbb{E}[\cdot]$ and $\sigma(\cdot)$ computed over a mini-batch. This is justified by our short horizon, so that $r_{f}$ is small relative to $R_{p}$, and by $r_{f}$ being approximately constant w.r.t. model parameters, hence having negligible impact on gradient directions. For simplicity, we ignore transaction costs and assume $R_{p}$ has finite second moments.
@@
VI-A7 Neural Wavelet Loss
Our low-/high-pass filters are implemented as 1D FIR convolutions. As discussed in Section V-A1, the convolution weights $W$ correspond to the impulse responses, whose discrete Fourier transform gives the frequency responses. Therefore, the rFFT of the filter parameters directly yields the spectrum on the grid
@@ VI-A7 Neural Wavelet Loss
$\mathcal{L}_{low}$ $=\sum_{\omega}(\omega/\pi)^{p}\,\lvert G_{\text{low}}(\omega)\rvert^{2},$
$\mathcal{L}_{high}$ $=\sum_{\omega}(1-\omega/\pi)^{p}\,\lvert G_{\text{high}}(\omega)\rvert^{2},$
$\mathcal{L}_{overlap}$ $=\lvert G_{\text{low}}\rvert^{2}\cdot\lvert G_{\text{high}}\rvert^{2},$
$\mathcal{L}_{parseval}$ $=\big(\lvert G_{\text{low}}\rvert^{2}+\lvert G_{\text{high}}\rvert^{2}-2\big)^{2}.$
@@
$\begin{gathered}\rho=\frac{\lvert G_{\text{high}}\rvert^{2}}{\lvert G_{\text{low}}\rvert^{2}+\varepsilon},\\ \mathcal{L}_{ratio}=\max\bigl(\rho-\rho_{\max},\,0\bigr)+\max\bigl(\rho_{\min}-\rho,\,0\bigr).\end{gathered}$ (20)
@@
and the overall training objective is
@@ VI-B Validation Metric and Model Selection
$R_{\text{val}}=\sum_{t\in\mathcal{D}_{\text{val}}}\tilde{w}_{t}r_{t},$
or, in the risk-adjusted setting, a Sharpe-like ratio computed from the sequence $\{\tilde{w}_{t}r_{t}\}$.
Crucially, $L_{\text{val}}$ and $R_{\text{val}}$ are not necessarily monotonic. The threshold-based decision rule in Eq. (9) can flip actions near boundaries with minimal change in $L_{\text{val}}$ yet a large change in $R_{\text{val}}$. Moreover, cross-entropy weights samples roughly uniformly, whereas $R_{\text{val}}$ is dominated by a small number of large moves: improving calibration on many small-return samples may reduce $L_{\text{val}}$ with little impact on $R_{\text{val}}$, while better capturing large profitable moves can slightly worsen $L_{\text{val}}$ but substantially improve $R_{\text{val}}$.
Therefore, we use soft-label cross-entropy as a differentiable surrogate for optimization, but perform early stopping and hyperparameter selection solely based on $R_{\text{val}}$. In practice, we select the checkpoint with the highest validation ROI, using $L_{\text{val}}$ only as a diagnostic for optimization stability.
VI-B1 Regression Loss vs. Soft-label Loss
@@
To remove this layer, we form a soft target using a scaled sigmoid,
$y_{t}=\sigma(k\,\hat{\ell}_{t}),$
where $k>0$ controls saturation. The network outputs a logit $p_{t}$ with $P_{t}=\sigma(p_{t})$ and is trained with soft-label cross-entropy. At inference, we convert $p_{t}$ to an executable position using the same squashing and risk-budget mapping described in Section III, so the optimized scalar becomes the trading signal up to a fixed monotone transform.
Soft labels offer three advantages: they penalize sign errors on large moves more than small deviations, learn a decision signal rather than a return forecast, and integrate naturally with the Sharpe-oriented regularizer in Section VI-A.
Table II shows Renewable Energy results with identical architecture and optimization. We disable the Sharpe regularizer and only change the supervised loss (MSE, MAE, or soft-label). The soft-label objective yields higher ROI and Sharpe than regression despite comparable error on $\ell_{t}$, supporting soft-label training for trading signals.
TABLE II: Comparison of regression losses and soft-label loss on Renewable Energy with 10 seeds. All models share the same WaveLSFormer architecture and optimization setup, the Sharpe regularizer is disabled for all rows to isolate the effect of the supervised loss.
@@
VI-B2 Tanh Loss vs. Soft-label Loss
We observed that optimizing Transformers for U.S. equity trading is highly sensitive to the loss design. When using the tanh-based objective in Eq. (15), training often stagnates because the tanh nonlinearity saturates, producing near-zero gradients and effectively “deadlocking” the optimizer. We diagnosed this behavior by monitoring the $\ell_{2}$ norm of parameter gradients, which frequently collapsed toward zero, and found that an overly small initial learning rate further increases the likelihood of getting trapped in poor local optima. However, the soft-label loss overcomes this drawback and produces a stable training process.
@@
VII Experiment
VII-A Experiment Setup
VII-A1 Data Split and Leakage Prevention
We use five years of hourly U.S. equity data from 29.10.2020 to 29.10.2025 and adopt a temporal 70%/10%/20% split for training, validation, and test, corresponding to 5292, 756, and 1512 time steps. Due to limited computational resources, hyperparameters are tuned manually.

### v3

@@ 5 Architecture
5.1 Neural Wavelet Filters
@@
5.1.1 FIR Convolution Kernels and Classical Discrete–time Filters
In our neural wavelet front-end, each learnable filter is implemented as a one-dimensional finite impulse response (FIR) convolution kernel. Let $x[n]$ denote a discrete-time input signal and let $\bm{\theta}=(\theta_{0},\dots,\theta_{L-1})$ be the convolution kernel of length $L$. A standard conv1d layer with stride 1 and no dilation computes
@@ 5.1.1 FIR Convolution Kernels and Classical Discrete–time Filters
Therefore, $e^{\mathrm{j}\omega n}$ is an eigenfunction of the system and the corresponding eigenvalue
@@
5.1.2 Differentiable Spectral Regularization via rFFT
Let $\bm{\theta}\in\mathbb{R}^{L}$ be a real-valued FIR kernel and $\hat{\bm{\theta}}=\mathcal{F}(\bm{\theta})$ its DFT. In matrix form, $\hat{\bm{\theta}}=\mathbf{F}\bm{\theta}$ with a fixed $\mathbf{F}\in\mathbb{C}^{L\times L}$. Hence $\mathcal{F}$ is linear and differentiable everywhere, with constant Jacobian $\mathbf{F}$. In practice we use rFFT, which is an optimized implementation of the same linear map specialized to real inputs.
@@ 5.1.2 Differentiable Spectral Regularization via rFFT
$\mathcal{L}_{spec}(\bm{\theta})=\sum_{k}\rho\!\left(|\hat{\theta}_{k}|\right),$ (12)
where $\rho(\cdot)$ is differentiable. By the chain rule, the gradient of spectral regularizer with respect to the time-domain kernel is
$\frac{\partial\mathcal{L}_{spec}}{\partial\bm{\theta}}=\mathbf{F}^{\mathsf{H}}\,\frac{\partial\mathcal{L}_{spec}}{\partial\hat{\bm{\theta}}}$,
where $\mathbf{F}^{\mathsf{H}}$ denotes the Hermitian transpose of $\mathbf{F}$, i.e. the inverse DFT up to a constant scaling factor. Equivalently, back-propagation through rFFT amounts to applying an inverse FFT to the frequency-domain gradient. Therefore, adding Eq. (12) on rFFT-transformed learnable kernels preserves differentiability and allows end-to-end training with standard optimizers.
To enforce complementary low/high-frequency coverage, we penalize high-frequency energy in the low-pass filter and low-frequency energy in the high-pass filter, add an overlap term to sharpen separation, and use Parseval and shape losses to constrain energy and frequency-response form; details are in Section 6.1.7.
5.1.3 Low-guided High-frequency Injection
@@
where $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{k}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ are projection matrices and $W_{O}$ is the output projection. The final fused representation is given by a gated residual injection:
@@
6 Training Design
Training choices, such as loss, optimizer, and learning-rate schedule, strongly influence trading outcomes. Because convenient objectives may misalign with deployment metrics, such as return, Sharpe and draw-down, low training loss can still generalize poorly. We therefore present: (i) a differentiable surrogate objective (soft labels from future log returns + Sharpe regularization), (ii) a stabilized optimization setup for heavy-tailed financial noise, and (iii) validation/checkpoint selection based on trading performance.
6.1 Loss Function Design
@@
6.1.1 Drawbacks of Regression Losses
Many financial time-series models train neural networks as return forecasters by minimizing point-wise regression losses such as MSE or MAE. Given a future log return $\ell_{t}$ as the ground truth, they train $f_{\theta}(\mathbf{X}_{t-1})$ via
$\mathcal{L}_{reg}=\mathbb{E}\big[loss(f_{\theta}(\mathbf{X}_{t-1}),\ell_{t})\big],\qquad loss\in\{\text{MSE},\ \text{MAE}\}.$ (14)
While statistically convenient, $\mathcal{L}_{reg}$ is misaligned with trading: performance is driven by the position $\tilde{w}_{t}$ and the strategy return profit $R_{t}=\tilde{w}_{t}\,f_{\theta}(\mathbf{X}_{t-1})$, not by the numeric forecast error of $f_{\theta}(\mathbf{X}_{t-1})$. Consequently, lower regression error does not necessarily translate to higher ROI/Sharpe. The mismatch is exacerbated by the asymmetric economic cost as a wrong sign under high exposure is far more damaging than a small magnitude error on a correct trade, whereas MSE/MAE penalize deviations symmetrically.
Moreover, under noisy and non-stationary returns, minimizing $\mathcal{L}_{reg}$ mainly encourages approximating $\mathbb{E}[f_{\theta}(\mathbf{X}_{t-1})]$, which offers no guarantee of improved $R_{t}$ or risk-adjusted performance. Therefore, we do not treat the network as a pure forecaster. Instead, we interpret its output as a continuous trading position $\tilde{w}_{t}\in[-1,1]$ (flat if $\tilde{w}_{t}\in[-0.01,0.01]$) and optimize a trading-aligned objective
$\mathcal{L}_{trade}=-\,\mathbb{E}\big[\tilde{w}_{t}\ell_{t}\big]+\lambda\,\Omega(\tilde{w}_{t}),$
@@ 6.1.1 Drawbacks of Regression Losses
6.1.2 Stock Direction and Tanh Position Loss
@@
6.1.3 Sigmoid Position Loss
@@
To emphasize economically significant moves, we weight each sample by the absolute log return $|\ell_{t}|$, i.e., $100\%$ and $50\%$ have equal $|\ell_{t}|$. The resulting Stock Sigmoid loss is
@@
6.1.4 Soft-Label Position Loss
@@
We therefore replace hard labels with probabilistic soft targets that reflect directional confidence $y_{t}=\sigma(k\,\ell_{t})$, where $k>0$ controls how quickly $y_{t}$ departs from $0.5$ as $|\ell_{t}|$ increases. We choose $k$ by calibration (e.g., enforcing $\sigma(k\,\ell_{t}^{(+5\%)})\approx 0.9$), which yields $k\approx 45$ in our setting; we fix $k=45$ for all experiments across industries for reproducibility. Given $y_{t}$ and $P_{t}$, we use the soft-label cross-entropy
@@
At inference time, we map the model output to a continuous position in [1,1][-1,1] using Eq. (9), where the sign encodes long, short and flat. The magnitude controls position size. Overall, soft labels provide smoother gradients, reduce over-emphasis on extreme returns, and align the learned signal with probabilistic directional confidence.
@@ 6.1.5 Overfitting Penalty
We cap the annualized ROI at $R_{\mathrm{ann}}^{\max}=1.0$ and treat it as a tunable risk-control parameter. With hourly data, $252$ trading days/year and $6.5$ hours/day, we have $H_{\mathrm{year}}=1638$. Assuming constant per-hour log return, the implied maximum batch log return is $L_{\mathcal{B}}^{\max}$, which yields the batch ROI threshold $T_{\mathcal{B}}$.
$L_{\mathcal{B}}^{\max}=\log(1+R_{\mathrm{ann}}^{\max})\,\frac{H_{\mathcal{B}}}{H_{\mathrm{year}}},\qquad T_{\mathcal{B}}=\exp(L_{\mathcal{B}}^{\max})-1=(1+R_{\mathrm{ann}}^{\max})^{H_{\mathcal{B}}/H_{\mathrm{year}}}-1$
We then penalize excess ROI using a one-sided quadratic hinge, where $\lambda_{\mathrm{roi}}$ controls the penalty strength.
@@
6.1.6 Sharpe Regularizer
@@
where $r_{f}$ is the risk-free rate. We omit subtle $r_{f}$ and use $S=\frac{\mathbb{E}[R_{p}]}{\sigma(R_{p})}$ with $\mathbb{E}[\cdot]$ and $\sigma(\cdot)$ computed over a mini-batch. This is justified by our short horizon, so that $r_{f}$ is small relative to $R_{p}$, and by $r_{f}$ being approximately constant w.r.t. model parameters, hence having negligible impact on gradient directions. For simplicity, we ignore transaction costs and assume $R_{p}$ has finite second moments.
@@
6.1.7 Neural Wavelet Loss
Our low-/high-pass filters are implemented as 1D FIR convolutions. As discussed in Section 5.1.1, the convolution weights $W$ correspond to the impulse responses, whose discrete Fourier transform gives the frequency responses. Therefore, the rFFT of the filter parameters directly yields the spectrum on the grid
@@ 6.1.7 Neural Wavelet Loss
$\mathcal{L}_{low}=\sum_{\omega}(\omega/\pi)^{p}\,\lvert G_{\text{low}}(\omega)\rvert^{2},\qquad\mathcal{L}_{high}=\sum_{\omega}(1-\omega/\pi)^{p}\,\lvert G_{\text{high}}(\omega)\rvert^{2},$
$\mathcal{L}_{overlap}=\lvert G_{\text{low}}\rvert^{2}\cdot\lvert G_{\text{high}}\rvert^{2},\qquad\mathcal{L}_{parseval}=\big(\lvert G_{\text{low}}\rvert^{2}+\lvert G_{\text{high}}\rvert^{2}-2\big)^{2}.$
@@
$\rho=\frac{\lvert G_{\text{high}}\rvert^{2}}{\lvert G_{\text{low}}\rvert^{2}+\varepsilon},\qquad\mathcal{L}_{ratio}=\max\bigl(\rho-\rho_{\max},\,0\bigr)+\max\bigl(\rho_{\min}-\rho,\,0\bigr).$ (20)
@@
6.2 Validation Metric and Model Selection
After the discussion above, the overall training objective is
@@ 6.2 Validation Metric and Model Selection
$R_{\text{val}}=\sum_{t\in\mathcal{D}_{\text{val}}}\tilde{w}_{t}r_{t}.$
In the risk-adjusted setting, we also compute a Sharpe-like ratio from the sequence $\{\tilde{w}_{t}r_{t}\}$ as an auxiliary validation metric.
Crucially, the validation loss $\mathcal{L}_{val}$ and $R_{\text{val}}$ are not necessarily monotonic. The threshold-based decision rule in Eq. (9) can flip actions near boundaries with minimal change in $\mathcal{L}_{val}$ yet a large change in $R_{\text{val}}$. Moreover, cross-entropy weights samples roughly uniformly, whereas $R_{\text{val}}$ is dominated by a small number of large moves: improving calibration on many small-return samples may reduce $\mathcal{L}_{val}$ with little impact on $R_{\text{val}}$, while better capturing large profitable moves can slightly worsen $\mathcal{L}_{val}$ but substantially improve $R_{\text{val}}$.
Therefore, we use soft-label cross-entropy as a differentiable surrogate for optimization, but perform early stopping and hyperparameter selection solely based on $R_{\text{val}}$. In practice, we select the checkpoint with the highest validation ROI, using $\mathcal{L}_{val}$ only as a diagnostic for optimization stability.
6.2.1 Regression Loss vs. Soft-label Loss
@@
To remove this layer, we form a soft target using a scaled sigmoid $y_{t}=\sigma(k\,\hat{\ell}_{t})$, where $k>0$ controls saturation. The network outputs a logit $p_{t}$ with $P_{t}=\sigma(p_{t})$ and is trained with soft-label cross-entropy. At inference, we convert $p_{t}$ to an executable position using the same squashing and risk-budget mapping described in Section 3, so the optimized scalar becomes the trading signal up to a fixed monotone transform.
Soft labels offer three advantages: they penalize sign errors on large moves more than small deviations, learn a decision signal rather than a return forecast, and integrate naturally with the Sharpe-oriented regularizer in Section 6.1.
Table 2 shows Renewable Energy results with identical architecture and optimization. We disable the Sharpe regularizer and only change the supervised loss (MSE, MAE, or Soft-Label). The soft-label objective yields higher ROI and Sharpe than regression despite comparable error on $\ell_{t}$, supporting soft-label training for trading signals.
Table 2: Comparison of regression losses and soft-label loss on Renewable Energy with 10 seeds. All models share the same WaveLSFormer architecture and optimization setup, the Sharpe regularizer is disabled for all rows to isolate the effect of the supervised loss.
@@
6.2.2 Tanh Loss vs. Soft-label Loss
We observed that optimizing Transformers for U.S. equity trading is highly sensitive to the loss design. When using the tanh-based objective in Eq. (15), training often stagnates because the tanh nonlinearity saturates, producing near-zero gradients and effectively deadlocking the optimizer. We diagnosed this behavior by monitoring the $L_{2}$ norm of parameter gradients, which frequently collapsed toward zero, and found that an overly small initial learning rate further increases the likelihood of getting trapped in poor local optima. However, the soft-label loss overcomes this drawback and produces a stable training process.
@@
7 Experiment
7.1 Experiment Setup
7.1.1 Data Split and Leakage Prevention
We use five years of hourly U.S. equity data from 29.10.2020 to 29.10.2025 and adopt a temporal $0.7:0.1:0.2$ split for training, validation, and test, corresponding to 5292, 756, and 1512 time steps. Due to limited computational resources, hyperparameters are tuned manually.